# Training and Evaluating with Human Label Variation: An Empirical Study

### v3

@@ Abstract
Human label variation (HLV) challenges the standard assumption that a labelled instance has a single ground truth, instead embracing the natural variation in human annotation to train and evaluate models. While various training methods and metrics for HLV have been proposed, it is still unclear which methods and metrics perform best in what settings. We propose new evaluation metrics for HLV leveraging fuzzy set theory. Since these new proposed metrics are differentiable, we then in turn experiment with employing these metrics as training objectives. We conduct an extensive study over 6 HLV datasets testing 14 training methods and 6 evaluation metrics. We find that training on either disaggregated annotations or soft labels performs best across metrics, outperforming training using the proposed training objectives with differentiable metrics. We also show that our proposed soft metric is more interpretable and correlates best with human preference.111Code is available at: https://github.com/kmkurn/train-eval-hlv
@@ 1 Introduction
Human label variation (HLV) challenges the standard assumption that an example has a single ground truth (Plank, 2022). Without this assumption, model training and evaluation are no longer straightforward. For example, standard metrics such as accuracy assume a single ground truth label exists for each instance. To address this issue, recent work (Pavlick and Kwiatkowski, 2019; Peterson et al., 2019; Nie, Zhou, and Bansal, 2020; Uma et al., 2020, 2021; Cui, 2023; Gajewska, 2023; Maity et al., 2023; Wan and Badillo-Urquiola, 2023, inter alia) has proposed various methods to train models and metrics to evaluate their performance in the HLV context. However, the lack of a systematic study means it is still unclear which methods and metrics perform best in what settings.
@@
In natural language processing (NLP), most existing metrics for HLV represent human judgements as probability distributions over classes and employ information-theoretic measures such as divergence. It has largely ignored the field of remote sensing that has dealt with issues in model evaluation against fuzzy ground truths (Foody, 1996; Binaghi et al., 1999; Lewis and Brown, 2001; Pontius and Cheuk, 2006; Pontius and Connors, 2006; Silván-Cárdenas and Wang, 2008; Gómez, Biging, and Montero, 2008). In remote sensing, the task is to classify patches of a satellite image of a region into categories such as land, water, and so forth. As an image patch often contains multiple categories, the field has developed evaluation metrics that can handle this variation in the ground truth labels. Taking inspiration from remote sensing research, we represent human judgement distributions as degrees of membership over fuzzy sets and generalise standard metrics such as accuracy into their soft versions using fuzzy set operations. To the best of our knowledge, we are the first to propose employing soft metrics for HLV in NLP.
Because these soft metrics are differentiable, we next explore the use of these soft metrics as training objectives. Through an extensive study testing 14 training methods and 6 evaluation metrics on 6 HLV datasets, we show that training on either disaggregated annotations or soft labels often achieves the best performance across metrics. We also conduct a meta-evaluation on the evaluation metrics. Albeit modest, we find that our proposed soft metric correlates best with human preference. However, this suggests that there is still substantial future work to develop better evaluation metrics for HLV.
@@
To summarise, we: (a) propose soft metrics to measure model performance in the context of HLV, taking inspiration from remote sensing research; (b) propose new training methods based on the soft evaluation metrics; and (c) conduct an extensive study covering 6 HLV datasets, 14 training methods and 6 evaluation metrics to understand the best training methods and evaluation metrics.
@@ 2 Related Work
While evaluating in the HLV context is a relatively new concept in NLP, evaluating against a fuzzy222The terms “fuzzy” and “crisp” are sometimes used in remote sensing literature to mean “soft” and “hard” in NLP literature. reference is a well-studied area in remote sensing research. Early work by Foody (1996) used a measure of distance that is equivalent to twice the Jenson-Shannon divergence (Lin, 1991). Binaghi et al. (1999) argued that entropy-based measures are sensible only if the reference is crisp, and proposed the use of fuzzy set theory to compute the soft version of standard evaluation metrics such as accuracy. A similar approach was also proposed by Harju and Mesaros (2023) for audio data. Key to this approach is the minimum function used to compute the class membership scores given a fuzzy reference and a fuzzy model output. Other proposed functions include product (Lewis and Brown, 2001), sum (Pontius and Connors, 2006), and a composite of such functions designed to ensure diagonality of the confusion matrix when the reference and the model output match perfectly (Pontius and Cheuk, 2006; Silván-Cárdenas and Wang, 2008). Approaches that take the cost of misclassifications into account have also been proposed (Gómez, Biging, and Montero, 2008). In this work, we take the fuzzy set approach by Binaghi et al. (1999) and apply it to text classification tasks.
@@ 3 Evaluation Metrics for HLV
We now discuss the evaluation metrics relevant for HLV data, first on multiclass metrics (Section 3.1), then multilabel metrics (Section 3.2), and finally the relationship between these metrics (Section 3.3). In these discussions, we will also introduce our use of soft metrics and explain how they relate to their hard variant counterparts. Table 1 provides a summary of the existing and proposed/new evaluation metrics.
@@ 3.1 Multiclass Metrics
This metric is novel to this work, for evaluation in the context of HLV. This is a special case of the soft micro F1 score proposed later in Section 3.2.
@@
By noting that , soft accuracy can be interpreted as the proportion of the judgement distribution that is correctly predicted. This interpretation is similar to the hard counterpart (i.e. proportion of examples that are correctly predicted) and arguably more intuitive than that of existing metrics because it is stated directly in terms of the problem of predicting human judgement distribution. Other similarities include symmetry (due to the symmetry of ) and boundedness (between 0 and 1), where the lower and the upper bounds are achieved if and only if and respectively for all . Note also that soft accuracy is reduced to the standard hard accuracy when all rows of both and are one-hot vectors (i.e. no label variation).
While functions other than have been used to compute soft accuracy (see Section 2), they lead to the soft accuracy being in the case where . This is problematic because intuitively, good HLV evaluation metrics should produce the highest score when the predicted judgement distribution perfectly matches the true counterpart. Therefore, we use the function in this work.
@@ 3.2 Multilabel Metrics
The (hard) micro F1 score is a special case of this soft counterpart when and (i.e. no label variation). Furthermore, the soft accuracy in Section 3.1 is a special case of this soft micro F1 score when the unity constraints hold. Therefore, the soft micro F1 score is a general evaluation metric for classification tasks.
@@ 3.3 Relationship between Metrics
To analyse the relationship between metrics, we compute Pearson correlation between each pair of multiclass metrics (excluding the macro ones) similar to prior work (Chicco and Jurman, 2020). For a given number classes, we sample matrices and and compute the value of each metric on these pairs of and . The Pearson correlations between pairs of metrics are then computed based on these data points. We set for computational reasons. We report the correlation coefficients in Table 2.
Table 2 shows that the metrics are generally weakly correlated with each other, with the exception of soft accuracy and PO-JSD where the two are consistently strongly correlated (). This strong correlation with an established HLV metric gives an empirical assurance that our proposed soft accuracy is a sensible metric. The table also shows that soft accuracy is only moderately correlated with (hard) accuracy in most cases, where this correlation gets weaker when the number of classes is large (), or either or is dense ( or ). This finding suggests that hard and soft accuracy capture different aspects when humans disagree. Furthermore, the table shows that entropy correlation is poorly correlated with all metrics, which suggests that it is an outlier HLV metric.
@@
Consider the same definitions of and used in Section 3. The soft accuracy and the PO-JSD between and satisfy the following inequality:
@@
This fact implies that soft accuracy penalises a model more heavily than PO-JSD, as illustrated in Figure 1. It shows that a skewed predicted judgement distribution of has a soft accuracy of , much less than the PO-JSD (). The high PO-JSD is potentially misleading as it is very close to the maximum possible value of 1.
@@ 4 Training Methods for HLV
We now discuss the training methods. In total, we experiment with 14 HLV training methods, as summarised in Table 3. The first method (ReL) is the simplest and uses standard cross-entropy loss as the training objective based on prior work (Sheng, Provost, and Ipeirotis, 2008). The next 8 methods (MV, AR, ARh, AL, SL, SLMV, AE, AEh) are from the Learning with Disagreements shared task (Leonardelli et al., 2023) which also use standard cross-entropy loss. The next 3 methods (JSD, SmF1, SMF1) use soft metrics that are differentiable as the training objective. The last 2 methods (LA-min, LA-max) aggregate the cross-entropy losses of all annotations for a given input. In detail:
@@
JSD, SmF1 (soft micro F1), and SMF1 (soft macro F1) train using JSD, soft accuracy for multiclass or soft micro F1 for multilabel tasks, and soft macro F1 as the objective respectively (see Section 3 for definitions).777We exclude entropy correlation because it doesn’t have a unique maximiser. These approaches leverage the differentiability of soft metrics, resulting in a single objective for both training and inference. By directly optimising the evaluation metric at training time, we expect the models to exhibit superior performance, especially when evaluated with the corresponding metric.
@@ 5 Evaluation of HLV Training Methods
We now conduct an extensive study covering 6 datasets, 14 training methods, and 6 evaluation metrics across 2 language models to understand the performance landscape of HLV data.
@@ 5.1 Experimental Setup
| MFRC (Trager et al., 2022) | Moral sentiment | EN | E | 18 | 8 | ✓ | ✓ |
@@
We use 5 public datasets with disaggregated annotations that have been used in previous HLV studies, four of which are in English and one in Arabic. Table 4 shows that the datasets cover both crowd (domain non-specialist) and expert (domain specialist) annotators, as well as the number of examples, average number of annotations, and the language of the dataset. For HS-Brexit, MD-Agreement, and ArMIS, we use the same train–test splits as Leonardelli et al. (2023). For both ChaosNLI and MFRC, we use 10-fold cross-validation with 10% of the training portion used as a development set. For ChaosNLI, we include only the SNLI and MNLI subsets (each has 1.5K examples) because of their standard NLI setup of one premise and one hypothesis.
@@
We also include a dataset called TAG which consists of English texts describing legal problems written by non-expert legal help seekers. We have access to this dataset because of our collaboration with Justice Connect,888https://justiceconnect.org.au an Australian public benevolent institution999As defined by the Australian government: https://www.acnc.gov.au/charity/charities/4a24f21a-38af-e811-a95e-000d3ad24c60/profile providing free legal assistance to laypeople facing legal problems. Each text in the dataset was annotated by an average of 5.5 practising lawyers, who each selected one or more out of the 33 areas of law101010Including a special label “Not a legal issue”. that applied to the problem. Thus, it is a multilabel classification task. Example areas of law include Neighbourhood disputes and Housing and residential tenancies. Average inter-annotator agreement () over the areas of law is , which is modest.111111We compute this agreement on random 10% selection of the data due to the high computational cost. This is because lawyers often have different interpretations of the same problem due to their different legal specialisations and years of experience. However, these different interpretations are all valid as human label variation because these registered lawyers are subject matter experts who play a crucial part in interpreting the law. The dataset has a total of 11K texts collected between July 2020 and early December 2023. We randomly split the dataset 8:1:1 for the training, development, and test sets respectively. While it cannot be released publicly because it is based on real-world confidential legal requests from help-seekers, we include the TAG dataset due to its importance for the meta-evaluation later in Section 6.
@@ 5.2 Results
We compute the difference between the performance of a single run of an HLV training method and the mean performance of MV. We then report the mean differences across runs. We present the results on ArMIS in Figure 2 (the only non-English dataset), HS-Brexit and MD-Agreement in Figure 3 (Learning with Disagreements shared task datasets), and the remaining datasets in Figure 4 (multiclass and multilabel datasets). We report performance of MV in Table 7 in the Appendix. We make the following observations.
@@
Second, ReL and SL substantially outperform MV on most datasets across models and evaluation metrics. Furthermore, ReL and SL are often among the top-performing methods, even when compared to JSD and SmF1 evaluated using PO-JSD and soft accuracy/micro F1 respectively. In other words, ReL and SL are often on par with (or outperform in some cases with LLaMA) the training method whose learning objective is the same as the evaluation metric, demonstrating the strength of ReL and SL.
Third, ReL and SL are generally competitive with each other. However, there are two exceptions to this trend. First, on ArMIS, SL outperforms ReL. Second, on both HS-Brexit and the SNLI portion of ChaosNLI, ReL outperforms SL with LLaMA. For the former finding, however, ArMIS is an outlier dataset: it is the only Arabic dataset, has the smallest number of examples and annotations per example, and all methods generally have higher variance on this dataset. The latter finding suggests that ReL is better suited to large language models than SL.

### v4

@@ Abstract
Human label variation (HLV) challenges the standard assumption that a labelled instance has a single ground truth, instead embracing the natural variation in human annotation to train and evaluate models. While various training methods and metrics for HLV have been proposed, it is still unclear which methods and metrics perform best in what settings. We propose new evaluation metrics for HLV leveraging fuzzy set theory. Since these new proposed metrics are differentiable, we then in turn experiment with employing these metrics as training objectives. We conduct an extensive study over 6 HLV datasets testing 14 training methods and 6 evaluation metrics. We find that training on either disaggregated annotations or soft labels performs best across metrics, outperforming training using the proposed training objectives with differentiable metrics. We also show that our proposed soft micro F1 score is one of the best metrics for HLV data.111Code is available at: https://github.com/kmkurn/train-eval-hlv
@@ 1 Introduction
Human label variation (HLV) challenges the standard assumption that an example has a single ground truth (Plank, 2022). With HLV, different human annotations for the same instance are viewed as useful signal rather than undesirable noise. In this context, model training and evaluation are no longer straightforward. For example, standard metrics such as accuracy assume that a single ground truth label exists for each instance. To address this issue, recent work (Pavlick and Kwiatkowski, 2019; Peterson et al., 2019; Nie, Zhou, and Bansal, 2020; Uma et al., 2020, 2021; Cui, 2023; Gajewska, 2023; Maity et al., 2023; Wan and Badillo-Urquiola, 2023, inter alia) has proposed various methods to train models and metrics to evaluate their performance in the HLV context. However, the lack of a systematic study means it is still unclear which methods and metrics perform best.
@@
In natural language processing (NLP), most existing metrics for HLV represent human judgements as probability distributions over classes and employ information-theoretic measures such as divergence. It has largely ignored the field of remote sensing that has dealt with issues in model evaluation against fuzzy ground truths (Foody, 1996; Binaghi et al., 1999; Lewis and Brown, 2001; Pontius and Cheuk, 2006; Pontius and Connors, 2006; Silván-Cárdenas and Wang, 2008; Gómez, Biging, and Montero, 2008). In remote sensing, the task is to classify patches of a satellite image of a region into categories such as land, water, and so forth. As an image patch often contains multiple categories, the field has developed evaluation metrics that can handle this variation in the ground truth labels.
Taking inspiration from remote sensing research, we represent human judgement distributions as degrees of membership over fuzzy sets and generalise standard metrics such as accuracy into their soft versions using fuzzy set operations. In contrast to information-theoretic measures, each of these soft metrics has an intuitive interpretation that corresponds to the standard counterpart. For example, soft accuracy is interpreted as the proportion of correctly predicted judgement distributions. To the best of our knowledge, we are the first to propose employing soft metrics for HLV in NLP.
As these soft metrics are differentiable, we next explore the use of each soft metric as the training objective. We perform an extensive study testing 14 training methods and 6 evaluation metrics on 6 HLV datasets across 2 pretrained models of different sizes to investigate which training methods are best for HLV data across the different metrics. This study is then followed by an empirical meta-evaluation of the evaluation metrics to understand which evaluation metrics are best for HLV data. To the best of our knowledge, we are the first to perform such an empirical meta-evaluation of HLV evaluation metrics.
@@
To summarise, we: (a) propose soft metrics to measure model performance in the context of HLV, taking inspiration from remote sensing research; (b) propose new training methods based on the soft evaluation metrics; and (c) conduct an extensive study covering 6 HLV datasets, 14 training methods, 2 pretrained models, and 6 evaluation metrics to understand the best training methods and evaluation metrics.
We find that two of the simplest methods for HLV training perform surprisingly well, outperforming more complex methods including training with each soft metric as the objective and attaining the best performance in most cases. The first method considers each annotation of an instance as a separate instance-label pair. Training is then performed on this disaggregated data where the same instance may occur multiple times with a different label each. The second method trains the model to predict the full label distribution induced by the instance’s annotations using the standard cross-entropy loss. This is in contrast to the standard model training where the prediction target is the mode of the label distribution.
For evaluation metrics, we find that an existing metric based on Jensen-Shannon divergence (JSD) and our proposed soft micro F1 score are two of the best HLV metrics. Our analysis shows that in single-label classification, the two metrics are highly correlated but we prove that soft accuracy (because micro F1 is equal to accuracy in this case) is upper-bounded by the JSD-based metric. As a result, we show that the JSD-based metric can give a score close to its maximum value to wrong predictions, which is potentially misleading, unlike our proposed soft accuracy. We thus recommend future work in HLV to report both metrics but focus on our proposed soft metrics when interpretability is important.
@@ 2 Related Work
While evaluating in the HLV context is a relatively new concept in NLP, evaluating against a fuzzy222The terms “fuzzy” and “crisp” are sometimes used in remote sensing literature to mean “soft” and “hard” in NLP literature. reference is a well-studied area in remote sensing research. Early work by Foody (1996) used a measure of distance that is equivalent to twice the Jensen-Shannon divergence (Lin, 1991). Binaghi et al. (1999) argued that entropy-based measures are sensible only if the reference is crisp, and proposed the use of fuzzy set theory to compute the soft version of standard evaluation metrics such as accuracy. A similar approach was also proposed by Harju and Mesaros (2023) for audio data. Key to this approach is the minimum function used to compute the class membership scores given a fuzzy reference and a fuzzy model output. Other proposed functions include product (Lewis and Brown, 2001), sum (Pontius and Connors, 2006), and a composite of such functions designed to ensure diagonality of the confusion matrix when the reference and the model output match perfectly (Pontius and Cheuk, 2006; Silván-Cárdenas and Wang, 2008). Approaches that take the cost of misclassifications into account have also been proposed (Gómez, Biging, and Montero, 2008). In this work, we take the fuzzy set approach by Binaghi et al. (1999) and apply it to text classification tasks.
Meta-evaluation of HLV evaluation metrics
There is little existing work on the meta-evaluation of HLV metrics. Most studies used information theory-based measures such as cross-entropy (Uma et al., 2020; Leonardelli et al., 2023, inter alia), presumably because of the standard probabilistic approach in modern NLP. Rizzi et al. (2024) proposed some theoretical properties that an ideal soft metric should satisfy and performed a theoretical meta-evaluation of existing HLV metrics for single-label classification. In contrast, we propose new soft metrics for both single- and multilabel classification and perform an empirical meta-evaluation of these metrics.
@@ 3 Evaluation Metrics for HLV
We now discuss the evaluation metrics relevant for HLV data, first on multiclass metrics (Section˜3.1), then multilabel metrics (Section˜3.2), and finally the relationship between these metrics (Section˜3.3). In these discussions, we will also introduce our use of soft metrics and explain how they relate to their hard variant counterparts. Table˜1 provides a summary of the existing and proposed/new evaluation metrics.
@@ 3.1 Multiclass Metrics
This metric is novel to this work, for evaluation in the context of HLV. This is a special case of the soft micro F1 score proposed later in Section˜3.2.
@@
Interpretation of soft accuracy
By noting that , soft accuracy can be interpreted as the proportion of the judgement distribution that is correctly predicted. For example, if soft accuracy is 70%, then the model predicts 70% of judgement distribution correctly. This interpretation is similar to the hard counterpart (i.e. proportion of examples that are correctly predicted) and arguably more intuitive than that of existing metrics because it is stated directly in terms of the problem of predicting human judgement distribution.
Besides interpretation, soft accuracy’s similarities to the hard counterpart include symmetry (due to the symmetry of ) and boundedness (between 0 and 1), where the lower and the upper bounds are achieved if and only if and respectively for all . Note also that soft accuracy is reduced to the standard hard accuracy when all rows of both and are one-hot vectors (i.e. no label variation).
While functions other than have been used to compute soft accuracy (see Section˜2), they lead to the soft accuracy being in the case where . This is problematic because intuitively, good HLV evaluation metrics should produce the highest score when the predicted judgement distribution perfectly matches the true counterpart. Therefore, we use the function in this work.
@@ 3.2 Multilabel Metrics
The (hard) micro F1 score is a special case of this soft counterpart when and (i.e. no label variation). Furthermore, the soft accuracy in Section˜3.1 is a special case of this soft micro F1 score when the unity constraints hold. Therefore, the soft micro F1 score is a general evaluation metric for classification tasks.
@@ 3.3 Relationship between Metrics
To analyse the relationship between metrics, we compute Pearson correlation between each pair of multiclass metrics (excluding the macro ones) similar to prior work (Chicco and Jurman, 2020). For a given number classes, we sample matrices and and compute the value of each metric on these pairs of and . The Pearson correlations between pairs of metrics are then computed based on these data points. We set for computational reasons. We report the correlation coefficients in Table˜2.
Table˜2 shows that the metrics are generally weakly correlated with each other, with the exception of soft accuracy and PO-JSD where the two are consistently strongly correlated (). This strong correlation with an established HLV metric gives an empirical assurance that our proposed soft accuracy is a sensible metric. The table also shows that soft accuracy is only moderately correlated with (hard) accuracy in most cases, where this correlation gets weaker when the number of classes is large (), or either or is dense ( or ). This finding suggests that hard and soft accuracy capture different aspects when humans disagree. Furthermore, the table shows that entropy correlation is poorly correlated with all metrics, which suggests that it is an outlier HLV metric.
@@
Consider the same definitions of and used in Section˜3. The soft accuracy and the PO-JSD between and satisfy the following inequality:
@@
This fact implies that soft accuracy penalises a model more heavily than PO-JSD, as illustrated in Figure˜1. It shows that a skewed predicted judgement distribution of has a soft accuracy of , much less than the PO-JSD (). The high PO-JSD is potentially misleading as it is very close to the maximum possible value of 1.
@@ 4 Training Methods for HLV
We now discuss the training methods. In total, we experiment with 14 HLV training methods, as summarised in Table˜3. The first method (ReL) is the simplest and uses standard cross-entropy loss as the training objective based on prior work (Sheng, Provost, and Ipeirotis, 2008). The next 8 methods (MV, AR, ARh, AL, SL, SLMV, AE, AEh) are from the Learning with Disagreements shared task (Leonardelli et al., 2023) which also use standard cross-entropy loss. The next 3 methods (JSD, SmF1, SMF1) use soft metrics that are differentiable as the training objective. The last 2 methods (LA-min, LA-max) aggregate the cross-entropy losses of all annotations for a given input. In detail:
@@
JSD, SmF1 (soft micro F1), and SMF1 (soft macro F1) train using JSD, soft accuracy for multiclass or soft micro F1 for multilabel tasks, and soft macro F1 as the objective respectively (see Section˜3 for definitions).777We exclude entropy correlation because it doesn’t have a unique maximiser. These approaches leverage the differentiability of soft metrics, resulting in a single objective for both training and inference. By directly optimising the evaluation metric at training time, we expect the models to exhibit superior performance, especially when evaluated with the corresponding metric.
@@ 5 Evaluation of HLV Training Methods
We now conduct an extensive study covering 6 datasets, 14 training methods, and 6 evaluation metrics across 2 pretrained models to understand the performance landscape of HLV data. Table˜4 shows an overview of the 6 datasets. These datasets are selected to represent diverse factors such as language, type of annotators, number of annotations for each example, number of classes, presence of annotator identity, and type of classification tasks (i.e. single- vs multilabel). By using these diverse datasets, consistent patterns that emerge across datasets are more likely to hold true in general rather than just in a specific type of datasets.
@@ 5.1 Experimental Setup
| MFRC (Trager et al., 2022) | Moral sentiment | EN | E | 18 | 8 | ✓ | ✓ | |
| TAG (private dataset) | Legal area | EN | E | 11 | 33 | ✓ | ✓ |
@@
We use 5 public datasets with disaggregated annotations that have been used in previous HLV studies, four of which are in English and one in Arabic. For HS-Brexit, MD-Agreement, and ArMIS, we use the same train–test splits as Leonardelli et al. (2023). For both ChaosNLI and MFRC, we use 10-fold cross-validation with 10% of the training portion used as a development set. For ChaosNLI, we include only the SNLI and MNLI subsets (each has 1.5K examples) because of their standard NLI setup of one premise and one hypothesis.
@@
We also include a private dataset called TAG which consists of English texts describing legal problems written by non-expert legal help seekers. Because it is based on real-world confidential legal requests from help-seekers, we are unable to distribute the dataset. However, we still include the TAG dataset due to its importance for the meta-evaluation later in Section˜6. We have access to this dataset because of our collaboration with Justice Connect,888https://justiceconnect.org.au an Australian public benevolent institution999As defined by the Australian government: https://www.acnc.gov.au/charity/charities/4a24f21a-38af-e811-a95e-000d3ad24c60/profile providing free legal assistance to laypeople facing legal problems. Each text in the dataset was annotated by an average of 5.5 practising lawyers, who each selected one or more out of the 33 areas of law101010Including a special label “Not a legal issue”. that applied to the problem. Thus, it is a multilabel classification task. Example areas of law include Neighbourhood disputes and Housing and residential tenancies. Average inter-annotator agreement () over the areas of law is , which is modest.111111We compute this agreement on random 10% selection of the data due to the high computational cost. This is because lawyers often have different interpretations of the same problem due to their different legal specialisations and years of experience. However, these different interpretations are all valid as human label variation because these registered lawyers are subject matter experts who play a crucial part in interpreting the law. The dataset has a total of 11K texts collected between July 2020 and early December 2023. We randomly split the dataset 8:1:1 for the training, development, and test sets respectively.
@@ 5.2 Results
We compute the difference between the performance of a single run of an HLV training method and the mean performance of MV. We then report the mean differences across runs. We present the results on ArMIS in Figure˜2 (the only non-English dataset), HS-Brexit and MD-Agreement in Figure˜3 (Learning with Disagreements shared task datasets), and the remaining datasets in Figure˜4 (multiclass and multilabel datasets). We report performance of MV in Table˜7 in the Appendix. We make the following observations.
@@
Second, both ReL and SL substantially outperform MV in 68 out of 78 settings (87.2% of all settings),151515Negative cases for ReL: all metrics except PO-JSD on ArMIS (5 settings), LLaMA on HS-Brexit and MD-Agreement with accuracy (2), LLaMA on MD-Agreement with macro F1 score (1), LLaMA on TAG with hard metrics (2). Negative cases for SL: TwHIN-BERT on ArMIS with hard metrics (2), LLaMA on HS-Brexit with all metrics except entropy correlation and soft macro F1 score (4), LLaMA on MD-Agreement and TAG with hard metrics (4). where a setting is a 3-tuple of a dataset, a model, and an evaluation metric.161616We consider the SNLI and the MNLI portions of ChaosNLI as separate datasets for this purpose. Furthermore, ReL or SL is among the top-performing methods in 89.7% of all settings,171717Negative cases: TwHIN-BERT on HS-Brexit with accuracy (1 setting); RoBERTa on MFRC with entropy correlation, PO-JSD, and soft macro F1 score (3); LLaMA on MFRC with macro F1 score, PO-JSD, and both micro and macro soft F1 scores (4). including when compared to JSD and SmF1 evaluated using PO-JSD and soft accuracy/micro F1 respectively. Concretely, ReL and SL are on par with (or outperform in some cases with LLaMA) the training method whose learning objective is the same as the evaluation metric in 23 out of 26 settings (88.5%),181818We consider only PO-JSD and soft accuracy/micro F1 evaluation metrics to get the total of 26 settings. Negative cases: RoBERTa on MFRC (2 settings), LLaMA on MFRC with PO-JSD (1). demonstrating the strength of ReL and SL.
Third, ReL and SL are generally competitive with each other. However, there are two exceptions to this trend. First, on ArMIS, SL outperforms ReL. Second, on both HS-Brexit and ChaosNLI, ReL outperforms SL with LLaMA. For the former finding, however, ArMIS is an outlier dataset: it is the only Arabic dataset, has the smallest number of examples and annotations per example, and all methods generally have higher variance on this dataset. The latter finding suggests that ReL is better suited to large language models than SL.